OpenStudy (anonymous):

Simplify. –22 + (–5) ∙ 6 A. –162 B. –52 C. 8 D. 162

OpenStudy (anonymous):

a

OpenStudy (anonymous):

OOP

OpenStudy (uri):

-22-5*6 = -27*6 =-162

OpenStudy (anonymous):

thank you ^-^